QUESTION 1

  1. Below you are given a partial computer output from a multiple regression analysis based on a sample of 16 observations.

    Coefficients

    Standard Error

    Constant

    12.924

    4.425

    x1

    -3.682

    2.630

    x2

    45.216

    12.560

    Analysis of Variance

    Source of

    Variation

    Degrees of Freedom

    Sum of
    Squares

    Mean
    Square

    F

    Regression

    4853

    2426.5

    Error

    485.3

    Carry out the test to determine if there is a relationship among the variables at the 5% level. The null hypothesis should

    a.

    be rejected.

    b.

    not be rejected.

    c.

    be revised to test for multicollinearity.

    d.

    test for individual significance instead.
Answer #1

F calculated = 2426.5/485.3 = 5

Regression df = 2

Error df = 16 - 2 - 1 = 13

Hence,

p - Value = P(F(2, 13) > 5) = 0.0245

Since this p - Value is less than 0.05, we reject Ho.

Option A is correct.